Alf can dig a trench in two hours, Bert takes three hours, Charlie takes 5 hours and Dave takes six hours. How long will they collectively take to dig the same trench if they all work simultaneously at their own speeds?
Answer by nerdybill(7384)

Alf can dig a trench in two hours, Bert takes three hours, Charlie takes 5 hours and Dave takes six hours. How long will they collectively take to dig the same trench if they all work simultaneously at their own speeds?
.
Let x = time (hours) it takes collectively
then
x(1/2 + 1/3 + 1/5 + 1/6) = 1
multiply both sides by common denominator 30:
x(15 + 10 + 6 + 5) = 30
x(36) = 30
x = 30/36
x = 0.8333 hours
or
x = 50 minutes